OpenStudy (anonymous):

1. Let f(x)=x^3-2x. Simplify the expression f(x+h)-f(x)/h 2. Let f(x)=3x+2 and g(x)=x/x+1 3. Let f(x)=2x^2. Simplify the expression f(x+h)-f(x)/h

OpenStudy (anonymous):

what you need is replace x by (x+h) x^3-2x =[((x+h)^3-2(x+h))-(x^3-2x)]/h =[(x^3+3x^2h+3xh^2+h^3)-(2x+2h)-(x^3-2x)]/h =[(x^3+3x^2h+3xh^2+h^3-2x-2h-x^3+2x)]/h =[3x^2h+3xh^2+h^3-2h)]/h =h[3x^2+3xh+h^2-2)]/h =3x^2+3xh+h^2-2) if you used the limit when h->0 the answer should be 3x^2-2 which is the derivative

OpenStudy (anonymous):

# 2 is not clear #3 as # 1